# captionUnknown document class (or package)

### v1

@@ I Introduction
Mixture-of-Experts (MoE) architectures have emerged as a central design choice for training large language models (LLMs) at scale. By integrating multiple specialized subnetworks, known as experts, MoE achieves substantial performance gains without a proportional increase in computational cost. A lightweight gating mechanism dynamically routes each input token to only a small subset of experts, enabling sparse activation and efficient parallelism during training. This sparsity enables MoE to scale significantly while keeping training costs manageable, leading to its adoption in state-of-the-art models such as Switch Transformer [1], Mixtral [2], and DeepSeek-V3 [3]. As a result, MoE has become a foundational architecture for training and scaling modern LLMs.
@@
Despite its efficiency during training, MoE inference remains resource-intensive, especially in terms of GPU memory, due to growing sizes of modern MoE models. Even with sparse activation, the cumulative memory footprint often exceeds the capacity of a single GPU. For example, Mixtral-8×7B [2] requires over 80 GB of GPU memory, far exceeding the 16–24 GB typically available on edge servers with commodity GPUs, such as those equipped with NVDIA RTX 4090 or A4000 GPUs commonly found in university labs, hospitals, small and medium-sized enterprises [4, 5]. As a result, the standard practice is to offload inference to cloud servers with large-scale GPU clusters. However, relying solely on centralized cloud services presents many limitations: prohibitively high expenditures of renting cloud GPU instances, high response latency from network round-trips, and growing concerns over data privacy and regulatory compliance.
@@
Why collaborative edge inference for MoE? In response, distributed deployments of LLMs including MoE models on GPU-equipped edge servers has emerged as a compelling alternative. Many organizations (e.g., universities and research labs) coordinate underutilized lab servers or desktops with modest GPU capabilities in different locations [5]. By doing so, edge-based inference offers significantly lower infrastructure costs and improved data locality. However, these methods still incur high MoE inference latency, as they lack fine-grained analysis of MoE model architectures and expert activation patterns. A few recent works have proposed strategies for edge MoE, but they either target acceleration on a single edge device [6] or overlook the system and request heterogeneity in edge scenarios [7, 8]: resource configurations, request volumes, and expert usages can all vary across distributed edge servers. To address this, we propose a collaborative MoE inference architecture, inspired by Expert Parallelism (EP), which distributes each layer of experts across GPUs. As illustrated in Fig. 1, inference is performed collaboratively across multiple devices, with each server hosting only a subset of the experts due to memory constraints. When a request arrives at Server 1 that activates an expert not present locally, such as Expert 8, the computation can still be completed by invoking the expert remotely from another server like Server 2 or Server 3. This setup leverages idle edge resources to form a cooperative inference layer, mitigating memory bottlenecks without relying on cloud execution.
@@
Importance and challenges of expert placement. Despite its advantages, collaborative MoE edge inference still faces challenges, particularly inter-GPU/server communication latency. Existing distributed MoE inference methods [9, 8, 10] including EP solutions designed for datacenters are unsuitable in edge scenarios due to two reasons: (i) edge servers vary in GPU memory, compute capacity, and network reliability, unlike datacenters with uniform hardware and high-speed interconnects; and (ii) redirecting requests across locations is costly, making it essential to process in proximity to data, a core principle of edge computing. These constraints call for strategic expert placement to reduce inference latency by maximizing local execution and minimizing remote calls. Additionally, expert activation patterns vary across servers due to user behavior and workload differences. While stable over short periods, longer-term shifts can degrade performance if placements remain fixed. Lightweight, periodical placement adjustment is needed to adapt layouts as workloads evolve, without complex migration protocols or frequent re-optimization.
@@ II-A Expert Activation Patterns
A core property of MoE models is their sparse activation: only a small subset of experts is activated per layer for each input. While expert activations may appear evenly distributed across general-purpose tasks, we make a key observation: for specific task types, the activation patterns become highly skewed and task-dependent. As illustrated in Fig. 2, when the Mixtral [2] model processes arithmetic tasks versus ASCII word recognition tasks from BIG-bench dataset [11], the expert utilization at the first layer varies significantly. Arithmetic tasks exhibit a strongly skewed activation pattern where Expert 6 dominates, while Expert 1 is rarely used. In contrast, ASCII word recognition predominantly activates Expert 3.
@@ II-B The Need of Collaborative Inference
As discussed above, one way to accommodate low-frequency experts is through memory offloading. For instance, MoE-Infinity [12] enables single-device MoE inference by dynamically loading rarely used experts from RAM to GPU memory, while frequently activated experts remain cached on the GPU. However, such methods are limited to a single-machine setup and cannot take advantage of the collective computing power available in edge environments. In practice, edge servers often exhibit heterogeneous workloads and resource capacities. For example, Mooncake’s open-source online trace dataset [13] shows that ToolAgent requests occur much more frequently than conversational queries. Moreover, the hardware heterogeneity among servers, such as differences in GPU availability or compute capability, leads to resource imbalance: some nodes remain underutilized while others are bottlenecked. To address this, a natural idea is to redistribute the load across servers.
One straightforward solution is to redirect requests to idle servers, i.e., reroute incoming requests to another server that performs inference locally using memory offloading. While this improves load balancing to some extent, our experiments show that it remains suboptimal. To highlight this, we compare three methods in Table I. The experiment uses the Mixtral-8×7B model deployed across three simulated edge servers. Each server processes a distinct type of request drawn from three datasets: arithmetic reasoning, ASCII word recognition, and abstract narrative understanding (from BIG-bench [11]). MoE-Infinity (w/ LB) represents the request-redirection baseline. In contrast, Naive Collaboration deploys experts randomly across the servers and enables distributed inference with remote expert calls. Despite its simplicity, the collaborative setup achieves significantly lower average latency, showing more balanced load across all servers.
@@ III-B Problem Formulation
Consider a system consisting of $N$ edge servers, where each server $n\in[N]$ is equipped with $G_{n}$ GPUs. The MoE model contains $L$ layers, each with a set of experts denoted by $E_{l}$ for layer $l\in[L]$. Let $x_{n}^{t}$ denote the $t$-th input batch received by server $n$ and $X_{n}$ denote the set of all input batches assigned to $n$. An expert placement strategy is denoted by $\mathcal{P}=\{z_{n,g}^{e}\}$, where $z_{n,g}^{e}\in\{0,1\}$ indicates whether expert $e\in\cup_{l=1}^{L}E_{l}$ is ($z_{n,g}^{e}=1$) placed on GPU $g$ of server $n$, or not ($z_{n,g}^{e}=0$). The objective is to minimize the total end-to-end inference latency across all input batches and servers, formalized as:
@@
Let $f_{n}^{l}(e)$ denote the empirical activation frequency of expert $e\in E_{l}$ on server $n$ at layer $l$. To determine whether a remote invocation occurs, we define the indicator:
@@
$\min_{\mathcal{P}}\quad\sum_{n=1}^{N}\sum_{l=1}^{L}\sum_{e\in E_{l}}f_{n}^{l}(e)\cdot\mathbf{1}_{\text{remote}}(n,e)$ (2)
@@
$\sum_{n=1}^{N}\sum_{g=1}^{G_{n}}z_{n,g}^{e}\geq 1,\quad\forall e\in\bigcup_{l=1}^{L}E_{l}$
$\sum_{l=1}^{L}\sum_{e\in E_{l}}z_{n,g}^{e}\cdot m_{e}\leq\texttt{mem}_{n,g},\quad\forall n,g$
@@ III-C1 Layer-wise Expert Count Allocation
$v_{n,l}=-\sum_{e\in E_{l}}p_{e}\log_{2}p_{e}\quad\text{where }p_{e}=\frac{f_{n}^{l}(e)}{\sum_{e^{\prime}\in E_{l}}f_{n}^{l}(e^{\prime})}.$
This Shannon entropy captures the variety of expert usage: it is low when a few experts dominate (low diversity) and high when activations are spread evenly. Then, each expert count is initialized as $N_{n,l}=\left\lfloor\frac{M_{n}}{m_{e}}\cdot\frac{v_{n,l}}{\sum_{l=1}^{L}v_{n,l}}\right\rfloor$, with server memory defined as $M_{n}=\sum_{g}\texttt{mem}_{n,g}$. This heuristic reflects the intuition that more diverse activation patterns require a broader set of experts to be co-located. To satisfy the expert-coverage constraint from the problem formulation, i.e., $\sum_{n}N_{n,l}=E_{l}$ for each layer, we adjust the initial allocation through expert rebalancing. If the total expert count in a layer $l$ is below $E_{l}$, we iteratively borrow experts from layers $l^{\prime}$ that are currently over-provisioned. Servers are prioritized in descending order of memory capacity to ensure feasibility. This process repeats until all layer-wise totals match their required expert counts.
@@
Input : Total experts per layer $E_{l}$, expert size $m_{e}$,
@@ 12 $N_{l,\text{total}}=\sum_{n=1}^{N}N_{n,l}$;
13 while $N_{l,\text{total}}<E_{l}$ do
14    $l^{\prime}=\arg\max_{l^{\prime}}\left(\sum_{n=1}^{N}N_{n,l^{\prime}}\right)$;
@@ 20          $N_{l,\text{total}}\leftarrow N_{l,\text{total}}+1$;
21          if $N_{l,\text{total}}==E_{l}$ then
@@
This follows from the asymptotic equipartition property [14]: the typical set covering $(1-\delta)$ of the probability mass has cardinality lower bounded by $2^{H(\mathbf{p})-\delta\log E}$. ∎
@@ III-C2 Expert-to-Server Assignment
Algorithm 2 selects the specific experts to place on each server, given the per-layer expert counts $N_{n,l}$ from Algorithm 1. Each server $n$ maintains a layer-wise preference list $\mathcal{P}_{n}^{l}$ over experts in $E_{l}$, sorted by empirical activation frequency $f_{n}^{l}(e)$. We initialize the expert set $\mathcal{A}_{n}$ by selecting the top-$N_{n,l}$ most frequently activated experts from $\mathcal{P}_{n}^{l}$ for each layer $l$. To ensure full expert coverage across the system (i.e., each expert $e\in E_{l}$ is placed on at least one server), we iteratively identify unassigned experts and reallocate them to servers holding redundant experts. For each unassigned expert $e\in\mathcal{U}_{l}$, we prioritize placement on servers with fewest duplicate assignments. Within each such server, we replace the least-frequently used duplicate with $e$. This procedure balances expert coverage and preference while respecting the memory-aware expert counts computed earlier.
@@ 5    $\mathcal{A}_{n}^{l}\leftarrow$ Top-$N_{n,l}$ experts from $\mathcal{P}_{n}^{l}$;
10 $\mathcal{U}_{l}\leftarrow\{e\in E_{l}\mid\sum_{n=1}^{N}\mathbf{1}\{e\in\mathcal{A}_{n}^{l}\}=0\}$;
@@ 20     Update $\mathcal{U}_{l}$;
$\sum_{l=1}^{L}\sum_{e\in E_{l}}f_{n}^{l}(e)\cdot(1-\mathbf{1}_{\text{remote}}(n,e))=\sum_{l=1}^{L}\sum_{e\in\mathcal{A}_{n}\cap E_{l}}f_{n}^{l}(e)$
@@
Let $f_{n}^{l}(e)\in[0,1]$ denote the normalized activation frequency of expert $e\in E_{l}$ on server $n$ at layer $l$. Let $B_{n}=\sum_{l=1}^{L}N_{n,l}$ be the total expert budget for server $n$ given by Algorithm 1. Define the local utility:
$U_{n}(S)=\sum_{l=1}^{L}\sum_{e\in S\cap E_{l}}f_{n}^{l}(e),$
for any subset $S\subseteq\bigcup_{l=1}^{L}E_{l}$ of size at most $B_{n}$. Then, the greedy assignment $\mathcal{A}_{n}$ returned by Algorithm 2 satisfies:
@@ IV Implementation and Experimental Results
Based on the open-source code of MoE-infinity [12], we implement DanceMoE. Building upon the MoE-Infinity framework, we have developed a multi-server MoE inference system. This flexible architecture allows us to accommodate various hardware configurations and optimize resource utilization to meet the demands of various deployment scenarios.
@@
BigBench
(10s Poisson
arrival)
Uniform 48.55 17.28 9.43 21.66 6.44 6.04 3.44 5.08
@@
MultiData
(20s Poisson
arrival)
Uniform 24.85 30.63 36.48 30.65 9.20 10.26 13.31 10.92

### v2

# Accelerating Edge Inference for Distributed MoE Models with Latency-Optimized Expert Placement

@@ I Introduction
Mixture-of-Experts (MoE) architectures have emerged as a central design choice for training large language models (LLMs) at scale. By integrating multiple specialized subnetworks, known as experts, MoE achieves substantial performance gains without a proportional increase in computational cost. A lightweight gating mechanism dynamically routes each input token to only a small subset of experts, enabling sparse activation and efficient parallelism during training. This sparsity enables MoE to scale significantly while keeping training costs manageable, leading to its adoption in state-of-the-art models such as Switch Transformer [fedus2022switch], Mixtral [jiang2024mixtral], and DeepSeek-V3 [liu2024deepseek]. As a result, MoE has become a foundational architecture for training and scaling modern LLMs.
@@
Despite its efficiency during training, MoE inference remains resource-intensive, especially in terms of GPU memory, due to growing sizes of modern MoE models. Even with sparse activation, the cumulative memory footprint often exceeds the capacity of a single GPU. For example, Mixtral-8×7B [jiang2024mixtral] requires over 80 GB of GPU memory, far exceeding the 16–24 GB typically available on edge servers with commodity GPUs, such as those equipped with NVDIA RTX 4090 or A4000 GPUs commonly found in university labs, hospitals, small and medium-sized enterprises [nvdia4090, gpuion]. As a result, the standard practice is to offload inference to cloud servers with large-scale GPU clusters. However, relying solely on centralized cloud services presents many limitations: prohibitively high expenditures of renting cloud GPU instances, high response latency from network round-trips, and growing concerns over data privacy and regulatory compliance.
@@
Why collaborative edge inference for MoE? In response, distributed deployments of LLMs including MoE models on GPU-equipped edge servers has emerged as a compelling alternative. Many organizations (e.g., universities and research labs) coordinate underutilized lab servers or desktops with modest GPU capabilities in different locations [gpuion]. By doing so, edge-based inference offers significantly lower infrastructure costs and improved data locality. However, these methods still incur high MoE inference latency, as they lack fine-grained analysis of MoE model architectures and expert activation patterns. A few recent works have proposed strategies for edge MoE, but they either target acceleration on a single edge device [edgemoe] or overlook the system and request heterogeneity in edge scenarios [fast-edge-moe, go2025moetuner]: resource configurations, request volumes, and expert usages can all vary across distributed edge servers. To address this, we propose a collaborative MoE inference architecture, inspired by Expert Parallelism (EP), which distributes each layer of experts across GPUs. As illustrated in Fig. 1, inference is performed collaboratively across multiple devices, with each server hosting only a subset of the experts due to memory constraints. When a request arrives at Server 1 that activates an expert not present locally, such as Expert 8, the computation can still be completed by invoking the expert remotely from another server like Server 2 or Server 3. This setup leverages idle edge resources to form a cooperative inference layer, mitigating memory bottlenecks without relying on cloud execution.
@@
Importance and challenges of expert placement. Despite its advantages, collaborative MoE edge inference still faces challenges, particularly inter-GPU/server communication latency. Existing distributed MoE inference methods [shoeybi2019megatron, go2025moetuner, deepseek2025eplb] including EP solutions designed for datacenters are unsuitable in edge scenarios due to two reasons: (i) edge servers vary in GPU memory, compute capacity, and network reliability, unlike datacenters with uniform hardware and high-speed interconnects; and (ii) redirecting requests across locations is costly, making it essential to process in proximity to data, a core principle of edge computing. These constraints call for strategic expert placement to reduce inference latency by maximizing local execution and minimizing remote calls. Additionally, expert activation patterns vary across servers due to user behavior and workload differences. While stable over short periods, longer-term shifts can degrade performance if placements remain fixed. Lightweight, periodical placement adjustment is needed to adapt layouts as workloads evolve, without complex migration protocols or frequent re-optimization.
@@ II-A Expert Activation Patterns
A core property of MoE models is their sparse activation: only a small subset of experts is activated per layer for each input. While expert activations may appear evenly distributed across general-purpose tasks, we make a key observation: for specific task types, the activation patterns become highly skewed and task-dependent. As illustrated in Fig. 2, when the Mixtral [jiang2024mixtral] model processes arithmetic tasks versus ASCII word recognition tasks from BIG-bench dataset [srivastava2023beyond], the expert utilization at the first layer varies significantly. Arithmetic tasks exhibit a strongly skewed activation pattern where Expert 0 dominates, while Expert 1 is rarely used. In contrast, ASCII word recognition predominantly activates Expert 3.
@@ II-B The Need of Collaborative Inference
As discussed above, one way to accommodate low-frequency experts is through memory offloading. For instance, MoE-Infinity [moe-infinity] enables single-device MoE inference by dynamically loading rarely used experts from RAM to GPU memory, while frequently activated experts remain cached on the GPU. However, such methods are limited to a single-machine setup and cannot take advantage of the collective computing power available in edge environments. In practice, edge servers often exhibit heterogeneous workloads and resource capacities. For example, Mooncake’s open-source online trace dataset [qin2025mooncake] shows that ToolAgent requests occur much more frequently than conversational queries. Moreover, the hardware heterogeneity among servers, such as differences in GPU availability or compute capability, leads to resource imbalance: some nodes remain underutilized while others are bottlenecked. To address this, a natural idea is to redistribute the load across servers.
One straightforward solution is to redirect requests to idle servers, i.e., reroute incoming requests to another server that performs inference locally using memory offloading. While this improves load balancing to some extent, our experiments show that it remains suboptimal. To highlight this, we compare three methods in Table I. The experiment uses the Mixtral-8×7B model deployed across three simulated edge servers. Each server processes a distinct type of request drawn from three datasets: arithmetic reasoning, ASCII word recognition, and abstract narrative understanding (from BIG-bench [srivastava2023beyond]). MoE-Infinity (w/ LB) represents the request-redirection baseline. In contrast, Naive Collaboration deploys experts randomly across the servers and enables distributed inference with remote expert calls. Despite its simplicity, the collaborative setup achieves significantly lower average latency, showing more balanced load across all servers.
@@ III-B Problem Formulation
Consider a system consisting of $N$ edge servers, where each server $n\in[N]$ is equipped with $G_{n}$ GPUs. The MoE model contains $L$ layers, each with a set of experts denoted by $\mathcal{E}_{l}$ for layer $l\in[L]$. In practice, we have $|\mathcal{E}_{l}|=E$, for each layer $l$. Let $x_{n}^{t}$ denote the $t$-th input batch received by server $n$ and $X_{n}$ denote the set of all input batches assigned to $n$. An expert placement strategy is denoted by $\mathcal{P}=\{z_{n,g}^{e}\}$, where $z_{n,g}^{e}\in\{0,1\}$ indicates whether expert $e\in\cup_{l=1}^{L}\mathcal{E}_{l}$ is ($z_{n,g}^{e}=1$) placed on GPU $g$ of server $n$, or not ($z_{n,g}^{e}=0$). The objective is to minimize the total end-to-end inference latency across all input batches and servers, formalized as:
@@
Let $f_{n}^{l}(e)$ denote the empirical activation frequency of expert $e\in\mathcal{E}_{l}$ on server $n$ at layer $l$. To determine whether a remote invocation occurs, we define the indicator:
@@
$\min_{\mathcal{P}}\quad\sum_{n=1}^{N}\sum_{l=1}^{L}\sum_{e\in\mathcal{E}_{l}}f_{n}^{l}(e)\cdot\mathbf{1}_{\text{remote}}(n,e)$ (2)
@@
$\sum_{n=1}^{N}\sum_{g=1}^{G_{n}}z_{n,g}^{e}\geq 1,\quad\forall e\in\bigcup_{l=1}^{L}\mathcal{E}_{l}$
$\sum_{l=1}^{L}\sum_{e\in\mathcal{E}_{l}}z_{n,g}^{e}\cdot m_{e}\leq\texttt{mem}_{n,g},\quad\forall n,g$
@@ III-C1 Layer-wise Expert Count Allocation
$v_{n,l}=-\sum_{e\in\mathcal{E}_{l}}p_{e}\log_{2}p_{e}\quad\text{where }p_{e}=\frac{f_{n}^{l}(e)}{\sum_{e^{\prime}\in\mathcal{E}_{l}}f_{n}^{l}(e^{\prime})}.$
This Shannon entropy captures the variety of expert usage: it is low when a few experts dominate (low diversity) and high when activations are spread evenly. Then, each expert count is initialized as $N_{n,l}=\left\lfloor\frac{M_{n}}{m_{e}}\cdot\frac{v_{n,l}}{\sum_{l=1}^{L}v_{n,l}}\right\rfloor$, with server memory defined as $M_{n}=\sum_{g}\texttt{mem}_{n,g}$. This heuristic reflects the intuition that more diverse activation patterns require a broader set of experts to be co-located. To satisfy the expert-coverage constraint from the problem formulation, i.e., $\sum_{n}N_{n,l}\geq E$ for each layer, we adjust the initial allocation through expert rebalancing. If the total expert count in a layer $l$ is below $E$, we iteratively borrow experts from layers $l^{\prime}$ that are currently over-provisioned. Servers are prioritized in descending order of memory capacity to ensure feasibility. This process repeats until all layer-wise totals match their required expert counts.
@@
Input : Total experts per layer $E$, expert size $m_{e}$,
@@ 12 $N_{l,\text{total}}=\sum_{n=1}^{N}N_{n,l}$;
13
14for $l=1$ to $L$ do
15 while $N_{l,\text{total}}<E$ do
@@ 22          $N_{l,\text{total}}\leftarrow N_{l,\text{total}}+1$;
23          if $N_{l,\text{total}}==E$ then
@@
This follows from the asymptotic equipartition property [Cover]: the typical set covering $(1-\delta)$ of the probability mass has cardinality lower bounded by $2^{H(\mathbf{p})-\delta\log E}$. ∎
@@ III-C2 Expert-to-Server Assignment
Algorithm 2 selects the specific experts to place on each server, given the per-layer expert counts $N_{n,l}$ from Algorithm 1. Each server $n$ maintains a layer-wise preference list $\mathcal{P}_{n}^{l}$ over experts in $\mathcal{E}_{l}$, sorted by empirical activation frequency $f_{n}^{l}(e)$. We initialize the expert set $\mathcal{A}_{n}$ by selecting the top-$N_{n,l}$ most frequently activated experts from $\mathcal{P}_{n}^{l}$ for each layer $l$. To ensure full expert coverage across the system (i.e., each expert $e\in\mathcal{E}_{l}$ is placed on at least one server), we iteratively identify unassigned experts and reallocate them to servers holding redundant experts. For each unassigned expert $e\in\mathcal{U}_{l}$, we prioritize placement on servers with fewest duplicate assignments. Within each such server, we replace the least-frequently used duplicate with $e$. This procedure balances expert coverage and preference while respecting the memory-aware expert counts computed earlier.
@@ 5    $\mathcal{A}_{n}^{l}\leftarrow$ Top-$N_{n,l}$ experts from $\mathcal{P}_{n}^{l}$;
10 $\mathcal{U}_{l}\leftarrow\{e\in\mathcal{E}_{l}\mid\sum_{n=1}^{N}\mathbf{1}\{e\in\mathcal{A}_{n}^{l}\}=0\}$;
@@ 20     Update $\mathcal{U}_{l}$;
$\sum_{l=1}^{L}\sum_{e\in\mathcal{E}_{l}}f_{n}^{l}(e)\cdot(1-\mathbf{1}_{\text{remote}}(n,e))=\sum_{l=1}^{L}\sum_{e\in\mathcal{A}_{n}\cap\mathcal{E}_{l}}f_{n}^{l}(e)$
@@
Let $f_{n}^{l}(e)\in[0,1]$ denote the normalized activation frequency of expert $e\in\mathcal{E}_{l}$ on server $n$ at layer $l$. Let $B_{n}=\sum_{l=1}^{L}N_{n,l}$ be the total expert budget for server $n$ given by Algorithm 1. Define the local utility:
$U_{n}(S)=\sum_{l=1}^{L}\sum_{e\in S\cap\mathcal{E}_{l}}f_{n}^{l}(e),$
for any subset $S\subseteq\bigcup_{l=1}^{L}\mathcal{E}_{l}$ of size at most $B_{n}$. Then, the greedy assignment $\mathcal{A}_{n}$ returned by Algorithm 2 satisfies:
@@ IV Implementation and Experimental Results
Based on the open-source code of MoE-infinity [moe-infinity], we implement DanceMoE. Building upon the MoE-Infinity framework, we have developed a multi-server MoE inference system. This flexible architecture allows us to accommodate various hardware configurations and optimize resource utilization to meet the demands of various deployment scenarios.
@@
BigBench (10s Poisson arrival) Uniform 48.55 17.28 9.43 21.66 6.44 6.04 3.44 5.08
@@
MultiData (20s Poisson arrival) Uniform 24.85 30.63 36.48 30.65 9.20 10.26 13.31 10.92